The store is selling a pearl necklace at $2500 with a markup of 65% jill has a coupon for 35 % off the original selling price. What is the total of the purchase price?
Found 2 solutions by rfer, god2012:
Answer by rfer(16322)

2500*0.65=$1625

Answer by god2012(113)

Solution:
Original selling price : $2500
Coupon Discount : 35%
Irrespective of the mark up , the necklace will be sold at its original selling price.
Therefore, Total discount = 2500 * 35 / 100 = $875
Purchase Price = 2500-875 = $1625